when buying a home theatre sustem for your local electronics store, you are offered one of two choices: pay $1,400 for the system outright, or oay $200 as an initial deposit and take a loan to pay off the redt over the jext year. the remainer of the ampunt may be paid off in monthlu insallments of $120. the lrevailing nominal interesr rate is 6%. how much would uou be laying for teh home thestre ststem if you chose the buy-on-credit option?
Answer #1

Total amount paid for the loan = $120 x 12 = $1,440

Total cost with the loan option ($) = Initial deposit + Total amount paid for the loan = 200 + 1,440 = 1,640

Therefore, if I choose buy-on-credit option, I pay $(1,640 - 1,400) = $240 extra.